OpenStudy (anonymous):

y = -3/2x + 4, –3x + 2y = 8: how would you graph these equations

OpenStudy (anonymous):

-3x+2y=8 2y=3x+8 y=3/2x+4 - eqn 2

OpenStudy (anonymous):